# Self-testing in the compiled setting via tilted-CHSH inequalities

### v1

@@ 4.2 Compiled self-test for tilted-CHSH inequalities
$\big{\|}(VX_{B}-(\sigma_{X}\otimes\mathbbm{1})V)|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$ $\leq\big{\|}V(X_{B}-\tilde{X}_{B})|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$+\big{\|}(V\tilde{X}_{B}-(\sigma_{X}\otimes\mathbbm{1})V)|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}.$
@@
$\big{\|}V(X_{B}-\tilde{X}_{B})|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}\leq\Bigg{(}\sum_{b\in\{0,1\}}\|(\tilde{X}_{B})^{b}P_{B}^{b}\|_{\mathrm{op}}^{2}\Bigg{)}\big{\|}(X_{B}-\tilde{X}_{B})|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
by the same calculation in 12, which after taking the expectation over $\mathsf{Enc}(x)=\chi$ and summing over $\alpha$, results in a contribution of $\delta_{4}$ when $x=0$ by 6, and $\delta_{4}+\delta_{\mathsf{QHE}}$ by Lemma 3.1 and 6 when $x=1$.
@@
Applying the triangle inequality, we end up with the sum of two terms. The first is given by
$\big{\|}(P_{B}^{0}\tilde{X}_{B}-\tilde{X}_{B}P_{B}^{1})|\Psi_{\alpha|\chi}\rangle\big{\|}^{2},$
which after taking the expectation over $\mathsf{Enc}(x)=\chi$ and summing over $\alpha$ results in a contribution of $\delta_{6}/4$ for $x=1$ by 11, and $\delta_{6}/4+\delta_{\mathsf{QHE}}$ for $x=0$ by Lemma 3.1 and 11. By the unitarity of $\tilde{X}_{B}$ the second is equal to $\big{\|}(P_{B}^{1}\tilde{X}_{B}-\tilde{X}_{B}P_{B}^{0})|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$. This is bounded by the same quantity as the previous, by noting that the proof of 11 is nearly identical with only the values $b$ of $P_{B}^{b}$ switched. ∎
@@
Let $\widetilde{\mathrm{Q}}$ be any compiled model which satisfies $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[S_{\theta,\phi}]\geq\eta_{\theta,\phi}^{\mathrm{Q}}-\epsilon$, $\widetilde{\mathrm{Q}}^{*}$ be the honest compiled model derived from Eq. 4.5, and let $V$ be the SWAP isometry defined in Eq. 4.14. Then there exists auxiliary states $\{|\mathsf{aux}_{\alpha|\chi}\rangle\}_{\alpha,\chi}$ such that
@@
where $\zeta_{x}=(\cos^{2}(\phi)\,\delta_{8,x}+\sin^{2}(\phi)\,\delta_{9,x})/4+(1-x)\cdot\delta_{7}+x\cdot\delta_{0}$, and $(\mathbb{C}^{2},\{|\varphi_{a|x}\rangle\},\{Q_{b|y}\})$ is the compiled counterpart of the bipartite model defined in Eq. 4.5.
@@
by 12 and 13. Finally, we note that Eqs. 4.15 and 4.17 imply the existence of states $\{|\mathsf{aux}_{\alpha|\chi}^{\prime}\rangle\}_{\alpha,\chi}$ such that
@@
$|\varphi_{a|x=0}\rangle$ =$=\begin{cases}\cos(\theta)|0\rangle_{B}\ \mathrm{if}\ a=0,\\ \sin(\theta)|1\rangle_{B}\ \mathrm{if}\ a=1,\end{cases}$
@@
$\big{\|}VN_{b|y}|\Psi_{\alpha|\chi}\rangle-$ $Q_{b|y}|\varphi_{\mathsf{Dec}(\alpha)|x}\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\big{\|}^{2}\leq\big{\|}VN_{b|y}|\Psi_{\alpha|\chi}\rangle-(Q_{b|y}\otimes\mathbbm{1})V|\Psi_{\alpha|\chi}\rangle\big{\|}^{2}$
$+\big{\|}(Q_{b|y}\otimes\mathbbm{1})V|\Psi_{\alpha|\chi}\rangle-Q_{b|y}|\varphi_{\mathsf{Dec}(\alpha)|\chi}\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\big{\|}^{2},$
which is bounded by Eqs. 4.20 and 4.21, after taking the expectation over $\mathsf{Enc}(x)=\chi$ and summing over $\alpha$, by $\zeta_{x}=(\cos^{2}(\phi)\,\delta_{8,x}+\sin^{2}(\phi)\,\delta_{9,x})/2+(1-x)\cdot\delta_{7}+x\cdot\delta_{0}$. ∎
@@
Eq. 4.19 shows that the generalized tilted-CHSH expressions are a self-test according to Eq. 3.21, completing the proof of Theorem 4.4. ∎

### v3

@@ 4.2 Compiled self-test for tilted-CHSH inequalities
$\big\|(VX_{B}-(\sigma_{X}\otimes\mathbbm{1})V)|\Psi_{\alpha|\chi}\rangle\big\|^{2}$ $\leq 2\big\|V(X_{B}-\tilde{X}_{B})|\Psi_{\alpha|\chi}\rangle\big\|^{2}$
$+2\big\|(V\tilde{X}_{B}-(\sigma_{X}\otimes\mathbbm{1})V)|\Psi_{\alpha|\chi}\rangle\big\|^{2}.$
@@
$2\big\|V(X_{B}-\tilde{X}_{B})|\Psi_{\alpha|\chi}\rangle\big\|^{2}\leq 2\Bigg(\sum_{b\in\{0,1\}}\|(\tilde{X}_{B})^{b}P_{B}^{b}\|_{\mathrm{op}}^{2}\Bigg)\big\|(X_{B}-\tilde{X}_{B})|\Psi_{\alpha|\chi}\rangle\big\|^{2},$
where we performed the same calculation in 12, which after taking the expectation over $\mathsf{Enc}(x)=\chi$ and summing over $\alpha$, results in a contribution of $2\delta_{4}$ when $x=0$ by 6, and $2(\delta_{4}+\mathrm{negl}(\lambda))$ by Lemma 3.1 and 6 when $x=1$.
@@
Taking the norm squared, we end up with the sum of two terms. The first is given by
$2\big\|(P_{B}^{0}\tilde{X}_{B}-\tilde{X}_{B}P_{B}^{1})|\Psi_{\alpha|\chi}\rangle\big\|^{2},$
which after taking the expectation over $\mathsf{Enc}(x)=\chi$ and summing over $\alpha$ results in a contribution of $\delta_{6}/2$ for $x=1$ by 11, and $2(\delta_{6}/4+\mathrm{negl}(\lambda))$ for $x=0$ by Lemma 3.1 and 11. By the unitarity of $\tilde{X}_{B}$ the second is equal to $\big\|(P_{B}^{1}\tilde{X}_{B}-\tilde{X}_{B}P_{B}^{0})|\Psi_{\alpha|\chi}\rangle\big\|^{2}$. This is bounded by the same quantity as the previous, by noting that the proof of 11 is nearly identical with only the values $b$ of $P_{B}^{b}$ switched. ∎
@@
Let $\widetilde{\mathrm{Q}}$ be any compiled model which satisfies $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[S_{\theta,\phi}]\geq\eta_{\theta,\phi}^{\mathrm{Q}}-\epsilon$, $\widetilde{\mathrm{Q}}^{*}$ be the compiled-counterpart of model given in Eq. 3.18, and let $V$ be the SWAP isometry defined in Eq. 4.9. Then there exists auxiliary states $\{|\mathsf{aux}_{\alpha|\chi}\rangle\}_{\alpha,\chi}$ such that for $x\in\{0,1\}$
@@
where $\zeta_{x}=(\cos^{2}(\phi)\,\delta_{8,x}+\sin^{2}(\phi)\,\delta_{9,x})/2+2(1-x)\cdot\delta_{7}+2x\cdot\delta_{0}$.
@@
by 12 and 13. Finally, we note that Eqs. 4.10 and 4.12 imply the existence of states $\{|\mathsf{aux}_{\alpha|\chi}^{\prime}\rangle\}_{\alpha,\chi}$ such that
@@
[…]
@@
$\big\|VN_{b|y}|\Psi_{\alpha|\chi}\rangle$- $Q_{b|y}|\varphi_{\mathsf{Dec}(\alpha)|x}\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\big\|^{2}\leq 2\big\|VN_{b|y}|\Psi_{\alpha|\chi}\rangle-(Q_{b|y}\otimes\mathbbm{1})V|\Psi_{\alpha|\chi}\rangle\big\|^{2}$
$+2\big\|(Q_{b|y}\otimes\mathbbm{1})V|\Psi_{\alpha|\chi}\rangle-Q_{b|y}|\varphi_{\mathsf{Dec}(\alpha)|\chi}\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\big\|^{2},$
which is bounded by Eqs. 4.15 and 4.16, after taking the expectation over $\mathsf{Enc}(x)=\chi$ and summing over $\alpha$, by $\zeta_{x}=(\cos^{2}(\phi)\,\delta_{8,x}+\sin^{2}(\phi)\,\delta_{9,x})+2(1-x)\cdot\delta_{7}+2x\cdot\delta_{0}$. ∎
@@
By noting that, for $x\in\{0,1\}$, the functions $\zeta_{x},\delta_{0},$ and $\delta_{7}$ are of the form $f(\epsilon)+\mathrm{negl}(\lambda)^{\prime\prime}$, where $f(\epsilon)\geq 0$ and $f(0)=0$, we see that Eq. 4.10 and Eq. 4.12 establish Eq. (4.2a), while Eq. 4.14 establishes Eq. (4.2b). Together this shows that the generalized tilted-CHSH expressions are a self-test according to Definition 4.5, completing the proof. ∎
Appendix A Appendix
A.1 Efficient quantum circuits and homomorphic encryption
To define a quantum homomorphic encryption scheme we require the following concepts from quantum cryptography.
Definition A.1.
A procedure $\mathcal{P}$ is quantum polynomial time (QPT) if:
  1. 1.
  2. 2.
A family of quantum states $\mathcal{F}$ is QPT (preparable) if there is a QPT $\mathcal{P}$ for preparing $\mathcal{F}$.
We now define a quantum homomorphic encryption (QHE) scheme. A formal definition of QHE first appeared in [BJ15]. We follow the description of QHE outlined in [KLVY23, Bra18]:
Definition A.2.
  1. (i)
  2. (ii)
  3. (iii)
    A QPT algorithm $\mathsf{Eval}$ which takes as input a classical description of a quantum circuit $\mathsf{C}:\mathcal{H}\otimes(\mathbb{C}^{2})^{\otimes n}\to(\mathbb{C}^{2})^{\otimes m}$ from $\mathcal{C}$, a quantum plaintext $|\Psi\rangle\in\mathcal{H}$ on a Hilbert space, a ciphertext $\chi$, and evaluates a quantum circuit $\mathsf{Eval}_{\mathsf{C}}(|\Psi\rangle\otimes|0\rangle^{\mathrm{poly}(\lambda,n)},\chi)$ producing a ciphertext $\alpha\in\{0,1\}^{\ell}$.
  4. (iv)
  1. 1.
    1. (a)
    2. (b)
  2. 2.
    (Security against efficient quantum distinguishers). Fix a secret key $\mathsf{sk}\leftarrow\mathsf{Gen}(1^{\lambda})$. Any quantum polynomial time adversary $\mathfrak{A}$ with access to $\mathsf{Enc}(\cdot,\mathsf{sk})$ (but does not know $\mathsf{sk}$) cannot distinguish between ciphertexts $\chi\leftarrow\mathsf{Enc}(x_{0},\mathsf{sk})$ and $\chi^{\prime}\leftarrow\mathsf{Enc}(x_{1},\mathsf{sk})$ with non-negligible probability in $\lambda$, where $x_{0}$ and $x_{1}$ are any plaintexts chosen by the adversary. That is
    for all pairs $(x_{0},x_{1})$.
The KLVY compilation procedure requires schemes that satisfy (1) and (2). QHE schemes satisfying (1) and (2) have been described in [Mah20, Bra18].
A.2 Additional proofs
Proof of Lemma 3.2.
Let $V_{y}^{(\lambda)}:\mathcal{H}^{{}^{\prime}(\lambda)}\to\mathbb{C}^{|\mathcal{B}|}\otimes\mathcal{H}^{{}^{\prime}(\lambda)}$ be the isometry defined by
$V_{y}^{(\lambda)}|\phi\rangle=\sum_{b\in\mathcal{B}}|b\rangle\otimes\sqrt{N_{b|y}^{{}^{\prime}(\lambda)}}|\phi\rangle,\ \forall|\phi\rangle\in\mathcal{H}^{{}^{\prime}(\lambda)}.$ (A.1)
Furthermore, let $U_{y}^{(\lambda)}$ be the unitary which satisfies $U_{y}^{(\lambda)}(|0\rangle\otimes|\phi\rangle)=V_{y}^{(\lambda)}|\phi\rangle$ for all $|\phi\rangle\in\mathcal{H}^{{}^{\prime}(\lambda)}$. Define the projectors,
$\tilde{N}_{b|y}^{(\lambda)}:=U_{y}^{(\lambda)\dagger}(|b\rangle\!\langle b|\otimes\mathbbm{1})U_{y}^{(\lambda)}.$ (A.2)
Since $|\mathcal{B}|$ is constant with respect to $\lambda$ and each $N_{b|y}^{{}^{\prime}(\lambda)}$ is QPT, the resulting PVMs $\{\tilde{N}_{b|y}^{(\lambda)}\}_{b\in\mathcal{B}}$ are QPT for every $y\in\mathcal{Y}$. For the sub-normalized states, let $|\tilde{\Psi}_{\alpha|\chi}\rangle\in\mathcal{H}^{{}^{\prime}(\lambda)}\otimes\tilde{\mathcal{H}}^{(\lambda)}$ be any purification [^3] Strictly speaking, since $\rho_{\alpha|\chi}^{(\lambda)}$ is sub-normalized, $|\tilde{\Psi}_{\alpha|\chi}^{(\lambda)}\rangle$ is equal to the purification of $\rho_{\alpha|\chi}^{(\lambda)}/{\mathrm{t}r}[\rho_{\alpha|\chi}^{(\lambda)}]$ weighted by ${\mathrm{t}r}[\rho_{\alpha|\chi}^{(\lambda)}]$, whenever ${\mathrm{t}r}[\rho_{\alpha|\chi}^{(\lambda)}]>0$. of $\rho_{\alpha|\chi}^{(\lambda)}$ with $\mathcal{H}^{{}^{\prime}(\lambda)}\cong\tilde{\mathcal{H}}^{(\lambda)}$, and define
$|\Psi_{\alpha|\chi}^{(\lambda)}\rangle:=|0\rangle\otimes|\tilde{\Psi}_{\alpha|\chi}^{(\lambda)}\rangle\in\mathbb{C}^{|\mathcal{X}|}\otimes\mathcal{H}^{{}^{\prime}(\lambda)}\otimes\tilde{\mathcal{H}}^{(\lambda)}=:\mathcal{H}^{(\lambda)}.$ (A.3)
Again, since $|\mathcal{X}|$ is constant with respect to $\lambda$ the (sub-normalized) states $|\Psi_{\alpha|\chi}^{(\lambda)}\rangle$ are QPT-preparable. Now, extend each $\tilde{N}_{b|y}^{(\lambda)}$ to act trivially on the purifying system $\tilde{\mathcal{H}}^{(\lambda)}$ by defining $N_{b|y}^{(\lambda)}:=\tilde{N}_{b|y}^{(\lambda)}\otimes\mathbbm{1}$. We observe
${\mathrm{t}r}\big[N_{b|y}^{(\lambda)}|\Psi_{\alpha|\chi}^{(\lambda)}\rangle\!\langle\Psi_{\alpha|\chi}^{(\lambda)}|\big]$ $={\mathrm{t}r}\Big[\tilde{N}_{b|y}^{(\lambda)}{\mathrm{t}r}_{\tilde{Q}}[|\Psi_{\alpha|\chi}^{(\lambda)}\rangle\!\langle\psi_{\alpha|\chi}^{(\lambda)}|]\Big]$ (A.4)
$={\mathrm{t}r}[\tilde{N}_{b|y}^{(\lambda)}(|0\rangle\!\langle 0|\otimes\rho_{\alpha|\chi}^{(\lambda)})]$
$={\mathrm{t}r}\big[\big(|b\rangle\!\langle b|\otimes\mathbbm{1}\big)U_{y}^{(\lambda)}\big(|0\rangle\!\langle 0|\otimes\rho_{\alpha|\chi}\big)U_{y}^{(\lambda)\dagger}\big]$
$={\mathrm{t}r}\Big[\sqrt{N_{b|y}^{{}^{\prime}(\lambda)}}\rho_{\alpha|\chi}^{(\lambda)}\sqrt{N_{b|y}^{{}^{\prime}(\lambda)}}\Big]=p^{(\lambda)}(\alpha,b|\chi,x),$
where $\tilde{Q}$ denotes the purifying system $\tilde{\mathcal{H}}^{(\lambda)}$. Since $\tilde{\mathcal{H}}^{(\lambda)}$ has the same dimensions as $\mathcal{H}^{(\lambda)}$, the PVMs $N_{b|y}^{(\lambda)}$ are indeed QPT. ∎
Proof of Theorem 3.3.
To begin, we write
$\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[P^{\dagger}P]$ $=\sum_{ij}\gamma_{i}^{*}\gamma_{j}\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[(A_{x})^{k_{i}}w_{i}(B_{0},B_{1})(A_{x})^{k_{j}}w_{j}(B_{0},B_{1})]$ (A.5)
$\sum_{ij:k_{i}\oplus k_{j}=0}\gamma_{i}^{*}\gamma_{j}\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[\bar{w}_{ij}(B_{0},B_{1})]$ (A.6)
$\approx_{\mathrm{negl}(\lambda)}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\Bigg(\sum_{ij:k_{i}\oplus k_{j}=0}\gamma_{i}^{*}\gamma_{j}\bar{w}_{ij}(B_{0},B_{1})\Bigg)|\Psi_{\alpha|\chi}\rangle$
$\sum_{ij}\gamma_{i}^{*}\gamma_{j}\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[(A_{x})^{k_{i}+k_{j}}\bar{w}_{ij}(B_{0},B_{1})]$ (A.7)
$\approx_{\mathrm{negl}(\lambda)}\sum_{ij:k_{i}\oplus k_{j}=0}\gamma_{i}^{*}\gamma_{j}\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\langle\Psi_{\alpha|\chi}|\bar{w}_{ij}(B_{0},B_{1})|\Psi_{\alpha|\chi}\rangle$
Where in the fifth line, we used the fact that Bob’s observables satisfy the canonical relations, so we can always replace the canonical monomial $\bar{w}_{ij}$ with $w_{i}w_{j}$. The final line is obtained by noting the square inside the expectation. We therefore conclude $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[P^{\dagger}P]\approx_{\mathrm{negl}(\lambda)}h$ for some $h\geq 0$, which implies $|\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[P^{\dagger}P]-h|\leq\mathrm{negl}(\lambda)$, and $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}[P^{\dagger}P]\geq h-\mathrm{negl}(\lambda)\geq-\mathrm{negl}(\lambda)$ as required. Lastly, it is straightforward to verify from the definition that for any Bell functional $I$ we have $\tilde{\mathop{\mathbb{E}}}_{\mathrm{S}}(I)$ recovers the expected value under $\widetilde{\mathcal{\mathrm{Q}}}$. ∎